# SIGGesture: Generalized Co-Speech Gesture Synthesis via Semantic Injection with Large-Scale Pre-Training Diffusion Models

### v1

@@ 4.5 Evaluation Results
Figure 4: The visualization results of the proposed method. The gestures in the red box are semantic gesture, which is labeled by the content in green box.
@@
User study. As we all known, the evaluation of generative task is very subjective. Although some metrics are introduced to evaluate the performance, there is a huge gap between metric evaluation and human visual perception. We conduct a user study to evaluate the performance of different baselines. Specifically, we randomly select 50 synthesized samples and shuffle results of different methods. Following previous works[52, 44], the participants are asked to scoring the gesture based on naturalness, rhythm appropriateness, diversity, and semantic consistence for the shuffled visual results.. It’s important to note that, for diversity, we enable users to score motion diversity under the condition of smooth and natural motion. The results of user study are shown in Table 3. The results show that the generated gestures of SIGGesture are dominantly preferred on four metrics over the comparing baselines. Especially, SIGGesture exceeds these baslines by a large margin. It can be concluded that large-scale pre-training benefits rhythm and diversity, while semantic injection is advantageous for semantic consistency. Overall, SIGGesture is capable of generating more realistic, synchronized, diverse, and understandable gestures that humans prefer.
@@
Table 3: The statistical data of user study by comparing the proposed method to baselines. The $\downarrow$" means the lower, the better. The $\uparrow$" means the higher, the better.
@@
DiffuseStyleGesture (IJCAI’23) [44] 0.78 0.82 0.65 0.57
TalkSHOW (CVPR’23)[48] 0.56 0.63 0.57 0.37
QPGesture (CVPR’23)[45] 0.62 0.72 0.53 0.56
EMAGE (24) [31] 0.80 0.81 0.79 0.60
SIGGesture 0.82 0.87 0.83 0.92
Quantitative Results. We compare our method with the baselines using four metrics on BEAT dataset. The comparison results are shown in Table 4. As discussed in recent works [43, 2], these metrics have obvious weakness because these value can not keep consistent with their visual results. For example, TalkSHOW[48] generates unnatural body movements based on the given audio, as shown in our supplementary video, but the results of these metrics can not show this gap. Besides, the SRGR can not show the difference of various methods, because SRGR only considers the ground truth semantic motion for a specific speech content. All the same, the proposed method can outperform previous works across most metrics. Although we only consider two modalities (audio and speaker identity), the proposed method can excels state-of-the-art model that employs all five modalities, such as DiffStyleGesture [44].
Table 4: The statistical data of quantitative results by comparing SIGGesture to the baselines. The $\downarrow$" means the lower, the better. The $\uparrow$" means the higher, the better.
[…]
DiffuseStyleGesture (IJCAI’23) [44] 10.14 0.233 0.504 11.975
TalkSHOW (CVPR’23)[48] 7.313 0.279 0.463 12.859
QPGesture (CVPR’23)[45] 19.921 0.209 0.453 9.438
EMAGE (24) [31] 5.430 0.272 0.679 13.075
@@
4.6 Ablation study
In this subsection, we evaluate the contribution of two main parts, large-scale pre-training and semantic injection, by user study. The comparison results are shown in Table 5. The findings reveal that the pre-training with noise data can improve the quality of basic rhythmic gesture. In addition to pre-training, the proposed semantic injection can significantly improve the quality of semantic gesture.
Table 5: The statistical results of ablation study. "w/o" means "without".
@@ 4.7 Generalization
To evaluate the generalizability, we test the proposed method on a broader set of audio samples, including out-of-domain English, Chinese and Japanese. The results are shown in Figure 5. It is also recommended to view the supplementary for more intuitive understanding and more examples. As shown in the visualization, the proposed method can accurately capture the key semantics and present precise gestures cross different language. The enhanced generalization capacity of our methodology is attributable to two principal components. Firstly, the employment of pseudo-labeled data can significantly improve the generalizability of rhythmic gestures. Secondly, disentangling the generation of rhythmic and semantic gestures by semantic injection can strengthen the controllability of semantic gesture synthesis. We have established an extensive collection of semantic gestures, encompassing a comprehensive range of gestural expressions. We leverage LLMs as a bridge to link the semantic gesture database and linguistic contexts, which has capable of robust generalization across diverse language. This method effectively addresses the challenges posed by the sparsity and long-tail distribution of semantic gestures. As a result, the proposed method has distinguished generalizability over different speech inputs.
Figure 5: The visual results of the proposed method on in-the-wild speeches (English, Chinese, and Japanese). The gestures in the red box are semantic gesture, which is labeled by the content in green box.
5 Conclusion
@@
In this paper, we present a novel method named SIGGesture for high-quality co-speech gesture synthesis. To achieve this, we propose a semantic injection method with LLMs to improve controllability, interpretability, and generalization ability. With the ability of LLMs, this method can be applied to other language without any adjustment. Besides, we propose a large-scale dataset from Internet videos for pre-training the diffusion-based synthesis model. With the usage of pre-training and fine-tuning paradigm, the generated gestures are more robust to the variation of the audio. The proposed datasets and related statistical experiments can benefit the research of co-speech gesture synthesis, including controllable gesture synthesis, gesture foundation model in the future.

### v2

@@ 4.4. Comparing Baselines
Figure 4. The visualization results of the proposed method. The gestures in the red box are semantic gesture, which is labeled by the content in green box.
4.5. Evaluation Results
4.5.1. Qualitative Results
For this task, there is lack of absolutely objective metric to evaluate the performance of various methods. Therefore, we strongly recommend the readers to view our demonstration video to gain an intuitive understanding of the qualitative outcomes. Figure 3 shows some synthesized results of these baselines conditioned on the same speech. From these results, we can observe that the synthesized co-speech gestures of the proposed method are more realistic, agile, and diverse than those of the baselines on BEAT datasets. The baselines suffer different extents of jittering in the results, especially lacking of semantic gesture and naturalness. DiffuseStyleGesture (Yang et al., 2023b) is also a diffusion-based method, which directly learns the origin motion representations. However,its results lack smoothness and naturalness (e.g. foot sliding). QPGesture (Yang et al., 2023a) exhibits noticeably slower and less varied motion, while ours demonstrates agility comparable to actual movement. For the specific semantic expressions, apart from our approach, other methods cannot effectively synthesize proper semantic gestures. More results are shown in Figure 4. The results indicate that our method can generate proper semantic gestures for the speech semantics. For example, ”free will” refers to the first row with open arms; ”left” and ”right” indicate pointing to the corresponding direction. These results show the surpassing capability of SIGGesture in synthesizing semantic accurate and natural gestures.
4.5.2. User study.
@@ 4.5.1. Qualitative Results
As we all known, the evaluation of generative task is very subjective. Although some metrics are introduced to evaluate the performance, there is a huge gap between metric evaluation and human visual perception. We conduct a user study to evaluate the performance of different baselines. Specifically, we randomly select 50 synthesized samples and shuffle results of different methods. Following previous works(Zhi et al., 2023; Yang et al., 2023b), the participants are asked to scoring the gesture based on naturalness, rhythm appropriateness, diversity, and semantic consistence for the shuffled visual results. It’s important to note that, for diversity, we enable users to score motion diversity under the condition of smoothness and naturalness. The results of user study are shown in Table 3. The results show that the generated gestures of SIGGesture are dominantly preferred on four metrics over the comparing baselines. Especially, SIGGesture exceeds these baselines by a large margin in semantic relevancy. It can be concluded that large-scale pre-training benefits rhythm and diversity, while semantic injection is advantageous for semantic consistency. Overall, SIGGesture is capable of generating more realistic, synchronized, diverse, and understandable gestures that humans prefer.
@@
Table 3. The statistical data of user study by comparing the proposed method to baselines. The $\downarrow$” means the lower, the better. The $\uparrow$” means the higher, the better.
Methods Nat.$\uparrow$ Rhy.$\uparrow$ Div.$\uparrow$ Sem.$\uparrow$
DiffuseStyleGesture 0.78 0.82 0.65 0.57
TalkSHOW 0.56 0.63 0.57 0.37
QPGesture 0.62 0.72 0.53 0.56
EMAGE 0.80 0.81 0.79 0.60
SIGGesture 0.82 0.87 0.83 0.92
Figure 5. The visual results of the proposed method on in-the-wild speeches (English, Chinese, and Japanese). The gestures in the red box are semantic gesture, which is labeled by the content in green box.
4.5.3. Quantitative Results.
We compare our method with the baselines using four metrics on BEAT dataset. The comparison results are shown in Table 4. As discussed in recent works (Tseng et al., 2023; Ao et al., 2023), these metrics have obvious weakness because these valued can not keep consistent with the visual results. For example, TalkSHOW (Yi et al., 2023) generates unnatural body movements based on the given audio, as shown in our supplementary video, but the results of these metrics can not show this gap. Besides, the SRGR can not show the difference of various methods, because SRGR only considers the ground truth semantic motion for a specific speech transcripts. Therefore, the SRGR of our method is lower than TalkSHOW. The reason is that this metric cannot accurately represent semantic performance. Because SRGR calculates the similarity (distance of skeleton) between the generated motion and the GT motion (with semantic labels and duration) in 3D space. For the same sentence, different semantic words can be expressed. All the same, the proposed method can outperform previous works across most metrics. Although we only consider two modalities (audio and speaker identity), the proposed method can excels state-of-the-art model that employs all five modalities, such as DiffStyleGesture.
Table 4. The statistical data of quantitative results by comparing SIGGesture to the baselines. The $\downarrow$” means the lower, the better. The $\uparrow$” means the higher, the better.
@@
DiffuseStyleGesture 10.14 0.233 0.504 11.975
TalkSHOW 7.313 0.279 0.463 12.859
QPGesture 19.921 0.209 0.453 9.438
EMAGE 5.430 0.272 0.679 13.075
@@
4.6. Ablation study
In this subsection, we evaluate the contribution of two main parts, large-scale pre-training and semantic injection, by user study. The comparison results are shown in Table 5. The findings reveal that the pre-training with noise data can improve the quality of basic rhythmic gesture. In addition to pre-training, the proposed semantic injection can significantly improve the quality of semantic gesture. The experimental results align with our hypothesis that pre-training with large-scale data can enhance the fundamental capabilities of synthesizing rhythmic gestures, and incorporating semantic injection with LLMs can accurately generate semantic gestures.
Table 5. The statistical results of ablation study. ”w/o” means ”without”.
@@ 4.7. Generalization
To evaluate the generalization ability, we test the proposed method on a broader set of audio samples, including out-of-domain English, Chinese and Japanese. The results are shown in Figure 5. It is also recommended to view the supplementary video for more intuitive understanding. As shown in the visualization, the proposed method can accurately capture the key semantics and present precise gestures cross different language. The enhanced generalization capacity of our methodology is attributable to two principal components. Firstly, the employment of pseudo-labeled data can significantly improve the generalization ability of rhythmic gestures. Secondly, disentangling the generation of rhythmic and semantic gestures by semantic injection can strengthen the controllability of semantic gesture synthesis. We have established an extensive collection of semantic gestures, encompassing a comprehensive range of gestural expressions. We leverage LLMs as a bridge to link the semantic gesture database with linguistic contexts, which enables robust generalization across diverse languages. This method effectively addresses the challenges posed by the sparsity and long-tail distribution of semantic gestures. As a result, the proposed method has distinguished generalization over different speech inputs.
5. Conclusion
@@
In this paper, we introduce a novel method called SIGGesture for high-quality co-speech gesture synthesis. Our approach leverages a semantic injection technique using LLMs to enhance controllability, interpretability, and generalization. Thanks to the capabilities of LLMs, our method can be seamlessly applied to other languages without requiring any complex adjustments. Additionally, we have developed a large-scale 3D gesture dataset sourced from Internet videos to pre-train our diffusion-based synthesis model. By employing a pre-training and fine-tuning paradigm, our generated gestures exhibit greater robustness to variations in audio input. The proposed datasets and accompanying statistical experiments aim to advance the field of co-speech gesture synthesis, including areas such as controllable gesture synthesis and gesture foundation models. Looking ahead, our research will focus on generating full-body animations with enhanced and detailed expressiveness.
6. Acknowledgements
We thank Wenhao Ge and Baocheng Zhang for their kindly support.
Appendix A Motion Representation
The diffusion model synthesizes the latent feature, and then the decoder of the Vector Quantized Variational Autoencoder (VQVAE) recovers the skeletal rotation from the latent space. VQVAE (Oord et al., 2017) has shown excellent performance in reconstructing complex content, including image, animation, video, etc. This section will present the details of the VQVAE model. We use high-quality skeletal rotations as motion representations and parameterize these rotations using rotation matrices relative to a reference T-pose. The original motion is denoted as =$\mathcal{M}=\{m_{0},m_{1},...,m_{L}\}$, ${m_{i}\in R^{J\times 9}}$, where ${J}$ is the number of skeletal bones and 9 is the dimension of rotation matrix. The initial FPS of different datasets is converted to 30 FPS for a unified standard. For a specific motion sequence $\mathcal{M}$, we use VQVAE to encode and quantize it into a finite codebook $\mathcal{Z}=\{z_{i}\}_{0}^{N}$, where ${N}$ is the size of codebook and ${z_{i}}$ is a specific lexeme. By using VQVAE, we can quantize the continuous motion into discrete tokens marked as $\mathcal{M}_{token}=\{\mathcal{T}_{0},...,\mathcal{T}_{L/4}\}$ with 4$\times$ down-sampling. The framework of VQVAE is shown in Figure 6. A piece of original motion $\mathcal{M}$ firstly is converted to a latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{L/4}\}$. Then, ${\hat{e}}=\{e_{0},e_{1},...,e_{L/4}\}$ is substituted by its nearest vector ${z_{j}}$ in the codebook, which can be denoted as $\arg\min\limits_{j}\parallel e_{i}-z_{j}\parallel$.
The encoding process can be represented by the following formula
(13) $\begin{split}\hat{e}_{i}=VQ_{encoder}(\mathcal{M}).\end{split}$
[…]
(14) ^$\begin{split}\hat{\mathcal{M}}=VQ_{decoder}(\hat{e}_{i}).\end{split}$
$VQ_{decoder}$ and $VQ_{decoder}$ are modelled by Convolutional Neural Networks. In the training process, the encoder, the decoder, and the codebook are optimized by the following loss function
(15) $\begin{split}\mathcal{L}_{VQ}=\mathcal{L}_{recon}+\parallel\hat{e}-sg(e)\parallel+\parallel sg(\hat{e})-e\parallel,\end{split}$
where $\mathcal{L}_{recon}$ is the reconstruction error and $sg[*]$ denotes the stop gradient operator. The reconstruction error is calculated by the following formula
[…]
where ${\mathcal{M}^{\prime}}$ and $\mathcal{M}^{\prime\prime}$ are the velocity and acceleration of the original motion sequence $\mathcal{M}$ respectively. All motions are encoded into latent space for training the diffusion model. In our pipeline, VQVAE serves as a fundamental model used to encode motions into latent space and to recover motions from latent space.
Figure 6. The architecture of VQVAE. It contains an encoder, a decoder and a codebook.
Appendix B User Study
We place significant importance on the effectiveness and fairness of our user study evaluations. To this end, we have recruited 30 volunteers (15 men and 15 women), including 5 researchers with expertise in gesture or animation synthesis. Each participant evaluates 250 co-speech gesture videos, which are generated from 50 audio clips using 5 different methods. To ensure that participants understand each metric and can accurately differentiate between them, we provide detailed instructions and training. This training covers the standard evaluation process and explains the differences between the various metrics. The detailed difference of these metric are as following.
The user study focus on the following aspects, Naturalness, Rhythm, Diversity, Semantic. Naturalness refers to how realistically and convincingly the generated motion movements mimic real-world behavior and physics. This includes evaluating whether the movements are fluid, smooth, and consistent with real movements and real-world physical laws. Rhythm metric typically evaluates how well the visual animation aligns with the rhythm or timing of the accompanying audio. This involves examining whether the gesture movements (mainly are hand movements) synchronize effectively with the beats or tempo of the audio. Diversity metric measures how much variety is present in the generated gesture animations produced in response to a specific audio. A higher score of diversity ensures that the synthesized co-speech gestures are not only synchronized with the audio but also rich and varied in a range of different movements and styles rather than repeating similar patterns or showing minimal variation. Semantic metric evaluates how well the animations convey the intended meaning, specific semantic or emotional content of the audio. A high Semantic score indicates that the generated gesture accurately interprets and represents the content of the audio.
Participants first browse the videos to familiarize themselves with the content. They then formally score the videos, with 5 animations from different methods displayed simultaneously on the screen. The differences in scores will reflect the performance variations among the different algorithms.
Appendix C Video-based Collected dataset
To ensure safety, raw data is collected from official TED and YiXi sources. Each video on TED and YiXi varies in terms of speaker’s gender, age, speech topics, and talking styles. This diversity enriches the dataset with a variety of gesture types. Finally, the dataset contains about 50% English and 50% Chinese. Following previous work (Yoon et al., 2019), the selection criteria for shots of interest include: containing speech content, having a stable background, displaying the entire upper body, facing the camera, and ensuring the body occupies at least 50% of the screen. These video clips are then processed through a gesture motion capture pipeline, depicted in Figure 7, which comprises five key components. Person detection utilizes Yolo-v8 (Jocher et al., 2023), while tracking employs the Sort algorithm (Bewley et al., 2016). (Bewley et al., 2016). Person key points are detected by using ViTPose (Xu et al., 2022), which are trained on a large-scale dataset. SMPL parameters fitting is like SMPLify (Bogo et al., 2016), which can boost the pose accuracy of SMPLX (Pavlakos et al., 2019). Post-processing part mainly contains motion smoothing and quality checking.
Figure 7. The process pipeline of video-based gesture motion capture.
Finally, the data is represented using the NEUTRAL SMPL-X skeleton. It took about 2 months and utilized 64 V100 GPUs for collecting the video-based gesture data.
Appendix D LLM prompt
The prompt is divided into three components: the task definition, a series of examples for few-shot learning, and the final task. During inference, the LLM is employed solely to generate accurate semantic gesture mappings from a predefined set. Our experiments utilize GPT-4 (Achiam et al., 2023) to perform this task and also involve a comparison of various prompting strategies, such as zero-shot and few-shot learning. In our findings, zero-shot learning prompts generally fail to produce appropriate semantic gestures. Conversely, few-shot learning prompts demonstrate effectiveness in performing the annotation task and yield meaningful results. It should be noted that these examples for few-shot learning contain both positive and negative examples. Below, we provide an example of the prompt, noting that some terms have been simplified due to page constraints.
Figure 8. An example of the LLM prompt for generating proper candidate semantic gesture.